OpenStudy (anonymous):

Find the distance between these points. A (5, 8), B(-3, 4) I got AB = 4√5 is this correct?

OpenStudy (jdoe0001):

yes \( \bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(5\quad ,&8)\quad &(-3\quad ,&4) \end{array}\qquad d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\\ \quad \\ \implies d = \sqrt{(-3-5)^2 + (4-8)^2}\implies d=\sqrt{(-8)^2+(-4)^2}\\ \quad \\ d=\sqrt{64+16}\implies d=\sqrt{80}\implies d=4\sqrt{5}\)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

I'm not sure why I'm having an issue with this one but it's not coming out right when I do it. Would you mind helping? Find the distance between these points. W(-6, -8), X(6, 8)

OpenStudy (anonymous):

I did d = \[\sqrt{(-6-6)^{2}+ (-8-8)^{2}}\] d=\[\sqrt{(-12)^{2} + (-16^{2}}\] d=\[\sqrt{144 + 256}\] d=400

OpenStudy (jdoe0001):

\(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(-6\quad ,&-8)\quad &(6\quad ,&8) \end{array}\qquad d = \sqrt{(x_2-x_1)^2 + (y_2-y_1)^2}\\ \quad \\ \implies d = \sqrt{(6-(-6))^2 + (8-(-8))^2}\implies d=\sqrt{(12)^2+(16)^2}\\ \quad \\ d=\sqrt{144+256}\implies d=\sqrt{400}\implies d=20\)

OpenStudy (anonymous):

Oh, duh! I forgot to find the square root. Blonde moment. Thank you so much! You're awesome!

OpenStudy (jdoe0001):

yw